How do I solve this problem by substitution?
2x+y=0
3x-2y=7

Answer by Alan3354(69443)

How do I solve this problem by substitution?
2x+y=0 --> y = -2x
3x-2y=7
------
Sub for y into 2nd eqn
2x - 2(-2x) = 7
7x = 7
x = 1
------
y = -2